# VL-Nav: Real-time Vision-Language Navigation with Spatial Reasoning

### v3

@@ Abstract
Vision-language navigation in unknown environments is crucial for mobile robots. In scenarios such as household assistance and rescue, mobile robots need to understand a human command, such as “find a person wearing black”. We present a novel vision-language navigation (VL-Nav) system that integrates efficient spatial reasoning on low-power robots. Unlike prior methods that rely on a single image-level feature similarity to guide a robot, our method integrates pixel-wise vision-language features with curiosity-driven exploration. This approach enables robust navigation to human-instructed instances across diverse environments. We deploy VL-Nav on a four-wheel mobile robot and evaluate its performance through comprehensive navigation tasks in both indoor and outdoor environments, spanning different scales and semantic complexities. Remarkably, VL-Nav operates at a real-time frequency of 30 Hz with a Jetson Orin NX, highlighting its ability to conduct efficient vision-language navigation. Results show that VL-Nav achieves an overall success rate of 86.3%, outperforming previous methods by 44.15%.
Figure 1: We propose VL-Nav, a real-time zero-shot vision-language navigation approach with spatial reasoning that integrates pixel-wise vision-language features and curiosity-based exploration for mobile robots. (a) Hallway: The wheeled robot is tasked with “find a man in gray” in a hallway. Unlike the classical frontier-based method (red line) and VLFM (green line), VL-Nav (blue line) leverages pixel-wise vision-language (VL) features from the “gray cloth” cue for spatial reasoning, selecting the most VL-correlated goal point and successfully locating the missing person. The value map shows that the “gray cloth” VL cue prioritizes the right-side area, marked by yellow square points. (b) Apartment: The robot is tasked with “Go to the tall white trash bin.” It detected two different-sized white trash bins in bottom camera observation. However, it assigns a higher confidence score (0.98) to the taller bin than the shorter one (0.48). These pixel-wise VL features are incorporated into the spatial distribution to select the correct goal point, guiding the robot toward the taller bin.
@@ I Introduction
Efficient navigation following human instructions in unseen environments is critical for autonomous robots, with various applications spanning from home assistance [30, 33, 34] to planetary exploration [27, 37, 29]. In such scenarios, robots are tasked with exploring and identifying targets instructed by humans, a challenge we define as vision-language navigation (VLN). Suppose a man wearing gray is reported missing and a command “find a man in gray” is given, the robot should not engage in random exploration. Instead, it should prioritize directions with visual features showing a stronger correlation to the language cues, such as detecting “gray cloth” within its field of view. Real-world VLN requires a navigation system that can: (1) interpret pixel-wise vision-language features, (2) adapt and perform robustly across various environments, and (3) operate in real-time on low-power platforms.
However, there are still no VLN systems that can fully address the three capabilities. Existing approaches can be broadly categorized into classical, end-to-end learning, and modular learning methods [13]. Classical approaches [9, 10, 31, 35], while efficient, struggle to integrate vision-language (VL) features. End-to-end learning methods [19, 26, 7, 28, 24] show promise but are computationally expensive, prone to overfitting in simulation, and exhibit poor generalization to out-of-distribution scenarios. Modular learning approaches [3, 23, 14, 5, 4], on the other hand, demonstrate robust real-world performance [13]. However, they often depend on extensive real-world robot training data and lack the capacity for human-like reasoning. The advent of vision-language models (VLMs) and large language models (LLMs) has further enhanced modular navigation approaches [12, 38, 16, 39]. For instance, Vision-Language Frontier Maps (VLFM) [39] utilized VLMs to extract language-grounded features directly from RGB images. This enables the creation of a semantic map, guiding exploration based on human semantic knowledge. However, their reliance on computationally intensive models limits their deployability on low-power platforms. Moreover, VLFM relies heavily on a single image-level feature similarity for goal selection, limiting its ability to exploit fine-grained VL cues.
To bridge this gap, we present vision-language navigation (VL-Nav), a novel navigation framework optimized for low-power robots, achieving zero-shot VLN at 30 Hz with an onboard computer. VL-Nav employs the spatial reasoning on both frontier-based and instance-based target points. VL-Nav initially generates frontier-based target points from the dynamic occupancy map using partial frontier detection. It confines the search to a manageable field of view, thereby reducing computational overhead. Additionally, instance-based target points are incorporated to emulate human search patterns, enabling the robot to approach and verify potential target objects, thereby enhancing the success rate. To select the most informative goal point, we introduce CVL spatial reasoning. This technique first transforms pixel-wise vision-language features into a spatial scoring distribution using a Gaussian mixture model. Each target point is then assigned a vision-language semantic score based on this distribution. Subsequently, a curiosity-driven weighting, which encourages the robot to explore unknown areas, is applied to adjust these scores into CVL scores. The target point with the highest CVL score is ultimately selected as the goal point. This CVL spatial reasoning process ensures that the selected goal point not only closely aligns with the human description but also guides the robot to explore unknown areas.
Once a goal is selected, VL-Nav utilizes a classic planner for real-time obstacle-aware path planning, enabling seamless adaptation to partially known environments. By combining pixel-wise vision-language features with curiosity-based exploration via the novel CVL spatial reasoning, VL-Nav surpasses all the baselines, achieving intelligent navigation while remaining computationally feasible for real-world deployment.
@@
  • We propose VL-Nav, an efficient VLN system optimized for low-power robots, achieving robust real-time performance at 30 Hz with an onboard computer.
  • We empower VL-Nav with spatial reasoning by integrating pixel-wise vision-language features and curiosity-driven exploration for more efficient VLN..
  • We conduct comprehensive evaluations across four real-world settings, showing that VL-Nav outperforms prior methods by 44.15% in diverse environments.
@@ II Related Works
End-to-end learning methods directly map sensory inputs to navigation actions using deep neural networks, typically trained with imitation learning (IL) or reinforcement learning (RL) losses [19, 26, 7, 28, 24]. Despite their ability to learn semantic priors for goal-directed exploration, end-to-end approaches suffer from significant limitations. These include high computational cost, reliance on large-scale training data, and a lack of interpretability. Furthermore, these methods often struggle to generalize beyond simulation due to domain gaps, as evidenced by significant performance drops when transitioning to real world [13].
Figure 2: An overview of the VL-Nav pipeline. VL-Nav processes inputs including prompts, RGB images, odometry poses, and LiDAR scans. The Vision-Language (VL) module conducts open-vocabulary pixel-wise detection to identify areas and objects related to the prompt, generating instance-based target points. Concurrently, the map module performs terrain analysis and manages a dynamic occupancy map. Frontier-based target points are then identified based on this occupancy map, along with the instance points, forming a candidate points pool. VL-Nav employs spatial reasoning to select the most effective goal point from this pool for path planning.
@@
Foundation Models for Navigation:
Recent advances in foundation models like VLMs and LLMs have further enhanced semantic navigation by incorporating natural language understanding into modular learning approaches [12, 38, 16, 39]. The Vision-Language Frontier Maps (VLFM) method [39] exemplifies this trend by leveraging a pre-trained VLM to extract semantic values directly from RGB images, enabling zero-shot semantic prior understanding for frontier-based exploration. VLFM achieves SOTA performance in the simulated environment, addressing the limitations of prior frontier-based and goal-oriented modular methods [35, 11, 3, 17]. However, VLFM’s reliance on computationally intensive VLMs restricts its deployability on resource-constrained platforms, and its focus on a single semantic similarity score for each observation does not adequately account for pixel-wise vision-language understanding and spatial reasoning, which is crucial for VLN. Their real-robot demonstration involves mounting a high-power-consumption RTX 4090 gaming laptop on the quadrupedal robot to handle computation-intensive modules, which is an impractical design for mobile-robot applications.
Strengths of VL-Nav:
VL-Nav significantly advances vision-language navigation (VLN) by addressing the limitations of classical, end-to-end, and modular learning methods. It employs a rolling occupancy grid and partial frontier detection, reducing computational overhead and enhancing adaptability in dynamic environments on low-power platforms. The integration of instance-based target points with spatial reasoning enables pixel-wise vision-language feature interpretation and context-aware goal selection. Unlike systems reliant on heavy computational models, VL-Nav optimizes for resource-constrained platforms and maintains real-time performance at 30 Hz. This makes VL-Nav not only practical for real-world applications but also superior in navigating complex and various environments efficiently.
@@ III Methodology
Figure 3: A brief illustration of VL Scoring. The pixel-wise open-vocabulary detection results are transferred into the spatial distibution via the Gaussian mixture model regularized by the FOV weighting (the gray arrows in the figure). Then the frontier-based and the instance-based target points will be assigned with VL score based on the distribution Equation (3).
This section presents our VL-Nav pipeline (Fig. 2), which leverages a rolling occupancy grid map, frontier-based and instance-based target points, and curiosity-vision-language (CVL) Scoring Policy to select and publish a navigation goal. This goal is not only aligned with the given prompt but also leads the robot to explore unknown areas more effectively.
III-A Rolling Occupancy Map: Occupancy Update
We represent the environment as a 2D occupancy grid $\mathcal{G}$, where each cell may be free $(0)$, unknown $(-1)$, or occupied $(100)$. Upon receiving new sensor data (a merged obstacle-and-terrain cloud $\mathcal{P}$), we update $\mathcal{G}$ using the following steps:
@@ III-A Rolling Occupancy Map: Occupancy Update
    Expand if needed. If any newly observed points lie outside the current map boundaries, expand the grid in place to preserve historical data.
@@
    Clear stale obstacles in forward FOV. For each cell marked occupied within the robot’s forward field of view (FOV), check whether it still corresponds to a point in $\mathcal{P}$. Any “stale” obstacle cells are re-labeled as free.
  3. 3.
    Inflate new obstacles. For each point in $\mathcal{P}$ within the sensor range $R$, mark the corresponding cell as occupied, along with a local neighborhood determined by an inflation radius $r_{\mathrm{inflation}}$.
  4. 4.
    Raycast unknown cells to free. Perform line-of-sight raycasting from the robot’s pose $(x_{r},y_{r},\theta_{r})$ across its angular range $\theta\in[\theta_{r}-\tfrac{\mathrm{hfov}}{2},\;\theta_{r}+\tfrac{\mathrm{hfov}}{2}]$. Any unknown cells $(-1)$ intersected by a ray are marked free $(0)$.
Algorithm 1 summarizes these main steps, returning the updated grid $\mathcal{G}$. This occupancy update leverages consistent line-of-sight logic and inflation to support safe navigation.
Algorithm 1 Rolling Occupancy Map Update (Key Steps)
0:  Grid $\mathcal{G}$, merged point cloud $\mathcal{P}$, robot pose $(x_{r},y_{r},\theta_{r})$, inflation radius $r_{\mathrm{inflation}}$, half-FOV $\mathrm{hfov}/2$, sensor range $R$.
1:  $\textsc{ExpandIfNeeded}(\mathcal{G},\mathcal{P})$
[…]
5:  return  Updated grid $\mathcal{G}$
III-B Frontier-based Target Points
@@
A cell $(m_{x},m_{y})$ is a frontier if it is free, $\mathcal{G}(m_{x},m_{y})=0$, and at least one neighbor is unknown, $\mathcal{G}(m_{x}+\Delta x,m_{y}+\Delta y)=-1$ for some $(\Delta x,\Delta y)$. After every grid update, only cells in the forward wedge are tested:
@@
$\begin{split}\Bigl{|}\mathrm{Angle}\bigl{(}(m_{x},m_{y}),(x_{r},y_{r})\bigr{)}-\theta_{r}\Bigr{|}\;&\leq\;\frac{\mathrm{hfov}}{2},\\ \text{and}\quad\|(m_{x},m_{y})-(x_{r},y_{r})\|\;&\leq\;R.\end{split}$
@@
III-C Instance-Based Target Points (IBTP)
A vision-language detector periodically reports candidate instance centers in the form $(q_{x},\;q_{y},\;\text{confidence})$, where $q_{x},q_{y}$ denote the estimated global coordinates for a potential target instance, and confidence quantifies how likely this detection is to match the desired instance. If confidence is larger than the detection threshold $\text{confidence is larger than the detection threshold }\tau_{\mathrm{det}}$, the candidate is retained; otherwise, it is discarded as too uncertain. Retained points will be downsampled via a voxel-grid filter if many candidates lie in close proximity.
This instance-based approach mimics human behavior when searching for an object: upon glimpsing something that might match the target, one naturally moves closer to confirm. VL-Nav likewise does not ignore intermediate detections. Instead, any candidate above a confidence threshold is treated as a valid goal candidate, allowing the robot to approach and verify whether it truly is an instance of interest. If the detection turns out to be incorrect, the robot continues exploring via frontiers or other instance cues, producing an accurate and robust strategy for zero-shot navigation toward a target instance.
III-D CVL Scoring Policy
Once frontier centroids and instance-based targets are gathered, the system computes an CVL score for each candidate goal $\mathbf{g}$. Let $(x_{r},y_{r})$ be the robot position, $\theta_{r}$ the heading, and $d(\mathbf{x}_{r},\mathbf{g})$ the Euclidean distance. The angular offset is

### v4

# VL-Nav: A Neuro-Symbolic Approach for Reasoning-based Vision-Language Navigation

@@ Abstract
Navigating unseen, large-scale environments based on complex and abstract human instructions remains a formidable challenge for autonomous mobile robots. Addressing this requires robots to infer implicit semantics and efficiently explore large-scale task spaces. However, existing methods, ranging from end-to-end learning to foundation model-based modular architectures, often lack the capability to decompose complex tasks or employ efficient exploration strategies, leading to robot aimless wandering or target recognition failures. To address these limitations, we propose VL-Nav, a neuro-symbolic (NeSy) vision-language navigation system. The proposed system intertwines neural reasoning with symbolic guidance through two core components: (1) a NeSy task planner that leverages a symbolic 3D scene graph and image memory system to enhance the vision-language models’ (VLMs) neural reasoning capabilities for task decomposition and replanning; and (2) a NeSy exploration system that couples neural semantic cues with the symbolic heuristic function to efficiently gather the task-related information while minimizing unnecessary repeat travel during exploration. Validated on the DARPA TIAMAT Challenge navigation tasks, our system achieved an 83.4% success rate (SR) in indoor environments and 75% in outdoor scenarios. VL-Nav achieved an 86.3% SR in real-world experiments, including a challenging 483-meter run. Finally, we validate the system with complex instructions in a 3D multi-floor scenario. Project Website: https://sairlab.org/vlnav/
@@ I Introduction
Efficient navigation following complex human instructions in completely unseen environments remains a formidable challenge for autonomous mobile robots. For example, consider a navigation task in the DARPA TIAMAT challenge [22]: “Today’s weather report indicates rain. Help Rob find an umbrella, an appropriate jacket, and shoes.” This marks a critical shift from naive command following vision-and-language navigaion [14] to reasoning-based vision-language navigation (VLN), where the robot must bridge massive logical gaps. The robot must first accurately understand the implicit semantics of the abstract instruction, reasoning out semantic inferences (e.g., “rain” implies waterproof gear) and performing object disambiguation to identify that the required targets are an umbrella, a rain jacket, and rain boots. This is essential to prevent the robot from looking for a normal jacket or sneakers as the target objects. Subsequently, it must efficiently explore the unknown environment to find all these required items within a limited task time.
Figure 1: Given the complex instruction, VL-Nav autonomously navigates a university building. Top: The generated 3D global map and robot trajectory (green line). The projected value map (rainbow squares on the floor) visualizes the NeSy scoring policy, which prioritizes task-relevant areas to guide efficient exploration. Bottom: Key reasoning milestones where the system (1) identifies stairs for floor transition, (2) grounds the target object using spatial relation (on black box), and (3) verifies the recipient based on visual attributes.
However, a significant gap remains in addressing such complex VLN. Classical semantic-free approaches [9, 10, 33] lack the linguistic reasoning necessary to comprehend abstract human instructions. Conversely, end-to-end learning methods, such as RL-based [19, 28, 7] and VLA-based approaches [38, 5, 37], are notoriously data-hungry and struggle with sim-to-real transfer. While recent foundation model-based modular architectures [36, 12, 35, 39] integrate advanced VLMs for zero-shot single-object navigation, their tight coupling of target verification with exploration constrains inferential capabilities, frequently resulting in the identification of incorrect items (e.g., random shoes instead of rain boots). Furthermore, most of these methods over-rely on neural semantic cues while neglecting the potential of geometric frontiers, leading to aimless wandering within explored regions. Consequently, existing approaches fundamentally lack two critical capabilities: (1) Accurate Task Reasoning to parse and decompose abstract multi-target instructions; and (2) Efficient Exploration Strategies to rapidly locate multiple targets in large-scale environments.
To seamlessly bridge this gap, we propose a NeSy vision-language navigation system that fundamentally intertwines neural reasoning with symbolic guidance. Specifically, our VL-Nav architecture consists of two main modules: (1) NeSy Task Planner: This planner logically decomposes abstract, multi-target instructions into atomic subtasks, dynamically issuing either exploration subtasks to gather information or go to subtasks to navigate to and report target objects. Crucially, the planner grounds VLM’s neural reasoning in a unified symbolic memory comprising a 3D scene graph and object-centric image memory, anchoring VLM decisions in precise spatial awareness and historical context. (2) NeSy Exploration System: Serving as the exploration backbone, this module couples neural semantic cues from lightweight VLMs with a symbolic heuristic function. This integration rapidly guides the robot through unknown environments to discover targets while minimizing unnecessary travel.
We comprehensively validate our proposed VL-Nav system in both simulations and extensive real-world robot environments. When evaluated in the navigation tasks in the DARPA TIAMAT challenge, our system demonstrated remarkable performance, achieving an 83.4% success rate (SR) in indoor settings and 75% in outdoor environments. Furthermore, real-world deployment across four diverse environments confirms the capability of generalizing from highly occluded, indoor domains to unstructured outdoor environments. Our real-world experiments achieved an 86.3% success rate, covering challenging scenarios such as a 483-meter long-range trajectory. Additionally, we showcased a complex multi-floor demonstration as shown in Fig. 1.
@@
  • We introduce VL-Nav, a NeSy system designed to resolve reasoning-based VLN tasks by intertwining neural semantic understanding with symbolic guidance.
  • We introduce a robust NeSy Task Planner to guide VLM reasoning and replanning within a unified symbolic memory. When coupled with our NeSy Exploration System, this approach facilitates rapid multi-target discovery and minimizes unnecessary travel.
  • VL-Nav achieves 83.4% and 75% SR in the indoor and outdoor settings on the navigation tasks of the DARPA TIAMAT Challenge, respectively. In real-world experiments, it attains an 86.3% SR, including long-range runs of up to 483 meters and demonstrations of complex tasks in multi-floor environments.
@@ II Related Works
End-to-end learning methods seek to directly map sensory inputs and instructions to navigation actions. Traditionally, these approaches relied on deep neural networks trained via imitation learning (IL) or reinforcement learning (RL) [19, 28, 7, 30, 26]. More recently, Vision-Language-Action (VLA) models [38, 5, 37] have emerged, attempting to unify vision, language, and continuous control into single massive transformer architectures. Despite their ability to inherently learn semantic priors for goal-directed exploration, these end-to-end approaches suffer from significant limitations when applied to real-world, multi-target VLN. They are notoriously data-hungry, computationally expensive to run in real-time, and lack interpretability. Most critically, they frequently fail to generalize beyond their training distributions, leading to severe performance degradation during sim2real transfer [13].
@@
Foundation Model-Based Modular Learning Methods:
Recent advances in foundation models like large Vision-Language Models (VLMs) and Large Language Models (LLMs) have further enhanced semantic navigation by incorporating natural language understanding into modular learning approaches [12, 35, 36, 39]. For instance, the Vision-Language Frontier Maps (VLFM) method [36] leverages a pre-trained VLM to extract semantic values directly from RGB images, enabling zero-shot semantic prior understanding for frontier-based exploration. Other methods like ApexNav [39] and SG-Nav [35] build open-vocabulary representations to find novel categories. However, these systems inherently operate under the premise of naive instruction following—where the “target” is explicitly named (e.g., “go to the L-shape sofa”). They lack the complex logic required for Reasoning-based VLN needed in abstract and multi-target tasks where explicit targets are not provided (e.g., inferring that “rain” requires an “rain jacket” instead of a random coat). Furthermore, executing heavy foundation models all the time during the navigation process bottlenecks real-time operation on resource-constrained platforms.
@@ III Methodology
Figure 2: System pipeline overview.Complex tasks are decomposed into subtasks via VLM reasoning, guided by the unified memory system. The first subtask always serves as input to the NeSy exploration system, which either guides exploration or sets a direct goal point for robot navigation.
To address the limitations of existing methods in resolving abstract and multi-target tasks, we introduce VL-Nav, a neuro-symbolic vision-language navigation system that intertwines neural semantic understanding with symbolic precision. As shown in Fig. 3, the architecture consists:
@@
    NeSy Task Planner logically decomposes complex, abstract instructions into atomic subtasks, exploration and go to, by guiding a VLM’s neural reasoning with a unified symbolic memory comprising a 3D scene graph and an object-centric image memory for task decomposition, replanning, and completion determination.
@@
    NeSy Exploration System couples neural semantic cues from lightweight open-vocabulary detection models (OVDMs) with a symbolic heuristic function, enabling efficient task-guided exploration of unknown environments without the computational bottlenecks of pure VLM-in-the-loop approaches.
III-A NeSy Task Planner
III-A1 Unified Memory System
The Unified Memory System consists of a 3D scene graph and an object-centric image memory. First, we construct a 3D scene graph comprising two types of nodes: object nodes and room nodes, inspired by Hydra [16]. Room node masks are generated using our room segmentation algorithm, which relies on morphological operations. To label these room nodes, we utilize a Large Language Model (LLM) that infers labels based on the object nodes connected to each room node. The edges between object and room nodes are established through spatial inclusion; specifically, an edge is created if an object’s centroid falls within a room’s segmentation mask.
Object nodes are generated by an open-vocabulary detector. Each object node stores four key attributes: (1) the object centroid; (2) the highest-confidence detection score; (3) the robot’s pose at the viewpoint of this detection; and (4) the corresponding best-viewpoint RGB image. This integration of the 3D scene graph with object-centric best-viewpoint image memory forms a unified symbolic memory, providing the spatial and relational grounding necessary for subsequent task decomposition and replanning.
III-A2 Task Decomposition and Replanning
Leveraging Qwen3-VL [1] as the VLM backbone, the planner decomposes complex task descriptions into atomic “exploration” and “go to” subtasks. The VLM maintains a dynamic target object list, querying the memory module to track and register potential candidates. Upon the completion of each subtask, a replanning phase will be triggered, generating updated subtasks based on the updated symbolic memory.
For target acquisition, the system employs a coarse-to-fine verification strategy. For instance, when locating a specific item like a “rain jacket,” the “go to” subtask is determined via a two-step process: (1) Symbolic Filtering: The 3D scene graph filters the available object nodes to propose the top-$k$ (e.g., top-3) candidates based on detection confidence. (2) Neural Verification: The VLM performs fine-grained reasoning on the saved best-view images and the neighbor scene graph nodes of these top candidates to determine which object best semantically matches the abstract instruction. Once the target is chosen, the planner publishes the saved best-view pose as the navigation goal, directing the exploration system to navigate to and report the identified target.
Figure 3: An overview of the NeSy exploration system. It processes inputs including prompts, RGB images, poses, and LiDAR scans. The Vision-Language (VL) module conducts open-vocabulary detection to identify objects related to the prompt, generating instance-based target points. Concurrently, the map module performs terrain analysis and manages a dynamic occupancy map. Frontier-based target points are then identified based on this occupancy map, along with the instance points, forming a candidate points pool. Finally, this system employs the NeSy scoring to select the most effective goal point.
Figure 4: A brief illustration of VL Scoring. The pixel-wise open-vocabulary detection results are transferred into the spatial distibution via the Gaussian mixture model regularized by the FoV weighting (the gray arrows). Then the frontier-based and the instance-based target points will be assigned with VL score based on the distribution Equation (3).
III-B NeSy Exploration System
Taking the subtask from the task planner, the NeSy Exploration System operationalizes high-level directives by fusing neural semantic cues from open-vocabulary detectors with symbolic geometric heuristics. The pipeline as shown in Fig. 3 proceeds from the generation of candidate goal points to the hybrid NeSy scoring policy, facilitating rapid, task-guided exploration and navigation in real-time.
III-B1 Frontier-based Target Points
@@
A cell $(m_{x},m_{y})$ is a frontier if it is free, $\mathcal{G}(m_{x},m_{y})=0$, and at least one neighbor is unknown, $\mathcal{G}(m_{x}+\Delta x,m_{y}+\Delta y)=-1$ for some $(\Delta x,\Delta y)$. After every grid update, only cells in the forward FoV wedge are tested:
@@
$\begin{split}\Bigl|\mathrm{Angle}\bigl((m_{x},m_{y}),(x_{r},y_{r})\bigr)-\theta_{r}\Bigr|\;&\leq\;\frac{\mathrm{hfov}}{2},\\ \text{and}\quad\|(m_{x},m_{y})-(x_{r},y_{r})\|\;&\leq\;R,\end{split}$
where $(x_{r},y_{r})$ and $\theta_{r}$ denote the robot’s current position and heading, $\mathrm{hfov}$ represents the horizontal field of view, and $R$ is the maximum effective sensor range.
@@
III-B2 Instance-Based Target Points (IBTP)
The vision-language detector shown in Fig. 3 VL Module periodically reports candidate instance centers in the form $(q_{x},\;q_{y},\;\text{confidence})$, where $q_{x},q_{y}$ denote the estimated global coordinates for a potential target instance, and confidence quantifies how likely this detection is to match the desired instance. If $\text{confidence is larger than the detection threshold }\tau_{\mathrm{det}}$, the candidate is retained; otherwise, it is discarded as too uncertain. Retained points will be downsampled via a voxel-grid filter if many candidates lie in close proximity.
This instance-based approach mimics human behavior when searching for an object: upon glimpsing something that might match the target, one naturally moves closer to confirm. VL-Nav likewise does not ignore intermediate detections. Instead, any candidate above a confidence threshold is treated as a valid goal candidate, allowing the robot to approach and verify whether it truly is an instance of interest. If the detection turns out to be incorrect, the robot continues exploring via frontiers or other instance cues.
III-B3 NeSy Scoring Policy
Once frontier-based and instance-based target points are gathered as candidate goals, the system computes the NeSy score for each candidate goal $\mathbf{g}$. Let $(x_{r},y_{r})$ be the robot position, $\theta_{r}$ the heading, and $d(\mathbf{x}_{r},\mathbf{g})$ the Euclidean distance. The angular offset is: